\begin{thebibliography}{10}

\bibitem{AdaFou} R.~Adams and J.~Fournier, {\it Sobolev Spaces}, Elsevier, San Diego, 2003.




%\bibitem{Che} X. Chen, Z. Chen and J. Wang, Heat kernel for nonlocal operators with variable-order. {\it Stoch. Proc. Appl.}, 130 (2020), 3574--3647.

% \bibitem{Cue} E. Cuesta, M. Kirane, A. Alsaedi and B. Ahmad, On the sub-diffusion fractional initial value problem with time variable order. {\it Adv. Nonlinear Anal.}, 10 (2021), 1301--1315.

\bibitem{Dan} R. Dang, Y. Cui, J. Qu, A. Yang, Y. Chen, Variable fractional modeling and vibration analysis
of variable-thickness viscoelastic circular plate. {\it Appl. Math. Model.} 110 (2022), 767--778.

\bibitem{DieFor} K. Diethelm and N. Ford, Analysis of fractional differential equations, {\it J. Math. Anal. Appl.}, 265 (2002), 229--248.

\bibitem{Eva} L.~Evans, {\em Partial Differential Equations}, Graduate Studies in Mathematics, V 19, American Mathematical Society, Rhode Island, 1998.

\bibitem{FanHu} W. Fan, X. Hu, and S. Zhu, Numerical reconstruction of a discontinuous diffusive coefficient in variable-order time-fractional subdiffusion,  {\it J. Sci. Comput.}, 96 (2023), 13.

%\bibitem{FanHu} W. Fan, X. Hu, and S. Zhu, Numerical reconstruction of a discontinuous diffusive coefficient in variable-order time-fractional subdiffusion,  {\it J. Sci. Comput.}, 96 (2023), 13.

%\bibitem{GanVon} L.~Gandossi and U.~Von Estorff, An overview of hydraulic fracturing and other formation stimulation technologies for shale gas production, Scientific and Technical Research Reports, Joint Research Centre of the European Commission; Publications Office of the European Union, 2015. doi:10.2790/379646.
%

%\bibitem{GorLuc} R.~Gorenflo, Y.~Luchko, and M.~Yamamoto,  Time-fractional diffusion equation in the fractional Sobolev spaces, {\it  Fract. Calc. Appl. Anal.}, 18 (2015), 799--820.

%\bibitem{GarGiu} R. Garrappa, A. Giusti, and F. Mainardi, Variable-order fractional calculus: A change of perspective, {\it Commun. Nonlinear Sci. Numer. Simul.}, 102 (2021), 105904.
\bibitem{GarGiu} R. Garrappa, A. Giusti, and F. Mainardi, Variable-order fractional calculus: A change of perspective, {\it Commun. Nonlinear Sci. Numer. Simul.}, 102 (2021), 105904.



\bibitem{Gor} R. Gorenflo, A. Kilbas, F. Mainardi and S. Rogosin, {\it Mittag-Leffler Functions, Related Topics and Applications}. Springer, New York, 2014.

\bibitem{HouLow} T. Hou, J. Lowengrub, and M. Shelley, Boundary integral methods for multicomponent fluids and multiphase materials, {\it J. Comput. Phys.}, 169 (2001),  302--362.

\bibitem{HouWu} T. Hou and X. Wu, A multiscale finite element method for elliptic problems in composite materials and porous media, {\it J. Comput. Phys.},  134 (1997),  169--189.

\bibitem{Jinbook} B. Jin, {\it Fractional differential equations--an approach via fractional derivatives}, Appl. Math. Sci. 206, Springer, Cham, 2021.

% \bibitem{JiaOu} L.~Jiang and N.~Ou, Bayesian inference using intermediate distribution based on coarse multiscale model for time fractional diffusion equations, {\it Multiscale Model. Simul.}, 16 (2018), 327--355.


%\bibitem{KiaSlo} Y. Kian, M. Slodi\v{c}ka, \'E. Soccorsi and K. Bockstal, On time-fractional partial differential equations of time-dependent piecewise constant order. arXiv:2402.03482.

 %\bibitem{JinMC} B.~Jin, B.~Li, and Z.~Zhou, Subdiffusion with a time-dependent coefficient: analysis and numerical solution, {\it Math. Comput.}, 88 (2019), 2157--2186.

%\bibitem{KiaSocYam} Y. Kian, E. Soccorsi and M. Yamamoto, On time-fractional diffusion equations with space-dependent variable order, {\it Ann. Henri Poincar\'e}, 19 (2018), 3855--3881.


%\bibitem{LeSty} K. Le and M. Stynes, An $\alpha$-robust semidiscrete finite element method for a Fokker-Planck initial-boundary value problem with variable-order fractional time derivative. {\it J. Sci. Comput.}, 86 (2021),  22.

%\bibitem{LiLuo} B. Li, H. Luo and X. Xie, Analysis of a time-stepping scheme for time fractional diffusion problems with nonsmooth data, {\it SIAM J. Numer. Anal.}, 57 (2019), 779--798.

%\bibitem{LiWanWan} B. Li, H. Wang, and J. Wang, Well-posedness and numerical approximation of a fractional diffusion equation with a nonlinear variable order, {\it ESAIM: M2AN}, 55 (2021), 171--207.

%\bibitem{LiWanZhe} Y. Li, H. Wang, X. Zheng, Analysis and simulation of optimal control for a two-time-scale fractional advection-diffusion-reaction equation with space-time dependent order and coefficients. {\it Multiscale Model. \& Simul.} 21 (2023), 1690--1716.

\bibitem{LiWan} Y. Li, H. Wang, and X. Zheng, Analysis of a fractional viscoelastic Euler-Bernoulli beam and
identification of its piecewise continuous polynomial order, {\it Fract. Calc. Appl. Anal.}, 26 (2023), 2337--2360.


\bibitem{LiaSty}  H. Liang and M. Stynes, A general collocation analysis for weakly singular Volterra integral equations with variable exponent, {\it IMA J. Numer. Anal.},  44 (2024), 2725--2751.



%\bibitem{LorHar} C. Lorenzo and T. Hartley, Variable order and distributed order fractional operators, {\it Nonlinear Dyn.}, 29 (2002), 57--98.


\bibitem{MclTho} W. McLean and V. Thom\'ee, Numerical solution of an evolution equation with a positive-type memory term. {\it J. Austral. Math. Soc. Ser. B} 35 (1993), 23--70.

%\bibitem{MeeSik} M.~Meerschaert and A.~Sikorskii, {\it Stochastic Models for Fractional Calculus}, De Gruyter Studies in Mathematics, 2011.

%\bibitem{MetKla} R.~Metzler and J.~Klafter, The random walk's guide to anomalous diffusion: A fractional dynamics approach, {\em Phys. Rep.}, 339 (2000), 1--77.

\bibitem{Mcl} W. McLean, V. Thom\'ee, L. Wahlbin,  Discretization with variable time steps of an evolution equation with a positive-type memory term. {\it J. Comput. Appl. Math.} 69 (1996), 49--69.

%\bibitem{Mcl07} W. McLean and K. Mustapha, A second-order accurate numerical method for a fractional wave equation. {\it Numer. Math.}, 105 (2007), 481--510.

%\bibitem{Mus} K. Mustapha and W. McLean, Discontinuous Galerkin method for an evolution equation with a memory term of positive type. {\it Math. Comp.}, 78 (2009), 1975--1995.

\bibitem{Pod} I. Podlubny, {\em Fractional Differential Equations}, Academic Press, 1999.

%\bibitem{PanPer} G. Pang, P. Perdikaris, W. Cai and G. Karniadakis, Discovering variable fractional orders of advection-dispersion equations from field data using multi-fidelity Bayesian optimization. {\it J. Comput. Phys.}, 348 (2017), 694--714.


\bibitem{SakYam} K.~Sakamoto and M.~Yamamoto, Initial value/boundary value problems for fractional diffusion-wave equations and applications to some inverse problems. {\em J. Math. Anal. Appl.} 382 (2011), 426--447.

%\bibitem{SonKar} F. Song and G. Karniadakis, Variable-order fractional models for wall-bounded turbulent flows. {\it Entropy}, 23 (2021), 782.



% \bibitem{SunZha}  H.~Sun, Y.~Zhang, W.~Chen and D.~Reeves, Use of a variable-index fractional-derivative model to capture transient dispersion in heterogeneous media, {\it J. Contam. Hydrol.}, 157 (2014) 47--58.

%\bibitem{SunChe} H. Sun, W. Chen and Y. Chen, Variable-order fractional differential operators in anomalous diffusion modeling. {\it Physica A}, 388 (2009), 4586--4592.

%\bibitem{SunCha} H. Sun, A. Chang, Y. Zhang and W. Chen, A review on variable-order fractional differential equations: mathematical foundations, physical models, numerical methods and applications. {\it Fract. Calc. Appl. Anal.}, 22 (2019), 27--59.

\bibitem{Tho} V.~Thom\'{e}e, {\em Galerkin Finite Element Methods for Parabolic Problems}, Lecture Notes in Mathematics 1054, Springer-Verlag, New York, 1984.



\bibitem{Da} D. Xu, Uniform $l^1$ behavior of the first-order interpolant quadrature scheme for some partial integro-differential equations, {\it Appl. Math. Lett.}, 117 (2021), 107097.
    
\bibitem{YiGuo}   L. Yi and B. Guo, An $h$-$p$ version of the continuous Petrov-Galerkin finite element method for Volterra integro-differential equations with smooth and nonsmooth kernels, {\it SIAM J. Numer. Anal.},  53 (2015), 2677--2704.
    
\bibitem{Zhe} X. Zheng, Two methods addressing variable-exponent fractional initial and boundary value problems and Abel integral equation, {\it CSIAM Trans. Appl. Math.}, to appear.

%\bibitem{JMAA} H. Wang and X. Zheng, Wellposedness and regularity of the variable-order time-fractional diffusion equations, {\it J. Math. Anal. Appl.}, 475 (2019), 1778--1802.

%\bibitem{Web}  J. Webb, Weakly singular Gronwall inequalities and applications to fractional differential equations, {\it J. Math. Anal. Appl.}, 471 (2019), 692--711.






\end{thebibliography}

\documentclass[final,1p,times]{elsarticle}
\usepackage{lmodern}
\usepackage{amssymb}
\usepackage{bm}
\usepackage{graphicx,graphics}
\usepackage{caption2}
\usepackage{psfrag}
\usepackage{amsmath,amsfonts}
\usepackage{booktabs}
\usepackage{amscd}
\usepackage{multirow}
\usepackage{lmodern}
\usepackage{float}
\usepackage{latexsym}
\usepackage{lscape}
\usepackage{color}

\newtheorem{theorem}{Theorem}[section]

\newcommand{\ds}{\displaystyle}

\newcommand{\om}{\Omega}
\newcommand{\p}{\partial}

\journal{}

\begin{document}

\bibliographystyle{plain}
\begin{frontmatter}

\title{Analysis of evolution equation with
variable-exponent memory modeling multiscale viscoelasticity
\footnote{Corresponding author: Xiangcheng Zheng}}

\author[1]{Yiqun Li}
\ead{YiqunLi24@outlook.com}

\author[2]{Xiangcheng Zheng}
\ead{xzheng@sdu.edu.cn}

\address[1]{School of Mathematics and Statistics, Wuhan University,
Wuhan 430072, China}

\address[2]{School of Mathematics, State Key Laboratory of Cryptography and Digital Economy Security, Shandong University, Jinan 250100, China}

\begin{abstract}
We investigate the well-posedness and solution regularity of an evolution equation with non-positive type variable-exponent memory, which describes multiscale viscoelasticity in materials with memory. The perturbation method is applied for model transformation, based on which the well-posedness is proved. Then the weighted solution regularity is derived, where the initial singularity is characterized by the initial value of variable exponent.
\end{abstract}

\begin{keyword}
  variable exponent, evolution equation, memory, well-posedness, regularity

\MSC 35R09

\end{keyword}

\end{frontmatter}

\section{Introduction}
This work considers the mathematical analysis of  the evolution equation with non-positive type variable-exponent memory, where the variable exponent  determined by the fractal dimension of the microstructures of viscoelastic materials describes the structure variations under the long-term cyclic loads, which in turn propagate to macro scales that eventually results in material failure \cite{Dan,FanHu,GarGiu,HouLow,HouWu,LiWan,LiaSty}
     \begin{equation}\label{VtFDEs}
     \begin{split}
         &~~~\p_t u(\bm x,t) -k(t)* \Delta u(\bm x,t) = f(\bm x,t), \quad (\bm x,t)\in \Omega\times (0,T];\\[0.05in]
         & \ds u(\bm x,0)=u_0(\bm x),\quad\bm x\in\Omega;\quad u(\bm x,t)=0,\quad (\bm x,t)\in \partial\Omega\times [0,T].
     \end{split}
     \end{equation}
    Here $\Omega\subset \mathbb R^d$ ($1\leq d \leq 3$) is a simply-connected bounded domain with the piecewise smooth boundary $\partial\Omega$ and convex corners, $T>0$, the source term $f(\bm x,t)$ and $u_0(\bm x)$ are given functions, and the convolution term is defined via the variable-exponent kernel $k(t)$ with $\alpha(t) \in (0,1)$
     \begin{equation}\label{k}
     \begin{split}
      k(t)*\Delta u(\bm x,t):=\int_0^tk(t-s)\Delta u(\bm x,s)ds,~~     k(t) : = \frac{t^{-\alpha(t)}}{\Gamma(1-\alpha(t))}.
     \end{split}
     \end{equation}

    The mathematical analysis for model \eqref{VtFDEs} with $\alpha(t) \equiv \hat \alpha$ for some $0<\hat \alpha <1$ has attracted extensive research activities, see e.g. \cite{MclTho,Da,YiGuo}, while, to our best knowledge, the variable-exponent  model \eqref{VtFDEs} remains untreated in the literature since the variable-exponent kernel $k(t)$ loses favorable properties  of its constant-exponent counterpart such as positive definiteness. Consequently, many established techniques for constant-exponent models do not apply.

    Recently, a perturbation method is developed in \cite{Zhe} to analyze the variable-exponent two-sided space-fractional problems, which splits the variable-exponent kernel \eqref{k} into its constant-order analogue and a low-order perturbation term. We adopt this approach to reformulate the model \eqref{ode1} to a more tractable form, based on which the well-posedness of the model \eqref{VtFDEs} and its regularity estimate are proved. Different from the resolvent estimate method in \cite{Zhe}, we employ special functions in solution expression for more subtle estimates. The developed results provide a theoretical support for numerical analysis and real applications of the proposed model. 

\section{Preliminaries}\label{sec2}
\subsection{Notations}
Let $L^p(\om)$ with $1 \le p \le \infty$ be the Banach space of $p$th power Lebesgue integrable functions on $\om$. For a positive integer $m$,
let  $ W^{m, p}(\Omega)$ be the Sobolev space of $L^p$ functions with $m$th weakly derivatives in $L^p(\om)$ (similarly defined with $\om$ replaced by an interval $\mathcal I$). Let  $H^m(\Omega) := W^{m,2}(\Omega)$ and $H^m_0(\Omega)$ be its subspace with the zero boundary condition up to order $m-1$. 
For a Banach space $\mathcal{X}$, let $W^{m, p}(0,T; \mathcal{X})$ be the space of functions in $W^{m, p}(0,T)$ with respect to $\|\cdot\|_{\mathcal {X}}$. All spaces are equipped with standard norms \cite{AdaFou}.

 Denote the eigenpairs of $-\Delta$ with Dirichlet boundary conditions by $\{\lambda_i,\phi_i\}_{i=1}^{\infty}$ in which $\{\phi_i\}_{i=1}^{\infty}$ form an orthonormal basis in $L^2(\om)$ and the eigenvalues $\{\lambda_i\}_{i=1}^\infty$ form a positive and  non-decreasing sequence \cite{Eva}.
  We introduce the Sobolev space $\check{H}^s(\Omega)$ for $s\geq 0$ by
$$ \check{H}^{s}(\Omega) := \big \{ q \in L^2(\Omega): \| q \|_{\check{H}^s(\Omega)}^2 : = \sum_{i=1}^{\infty} \lambda_i^{s} (q,\phi_i)^2 < \infty \big \},$$
which is a subspace of $H^s(\Omega)$ satisfying $\check{H}^0(\Omega) = L^2(\Omega)$ and $\check{H}^2(\Omega) = H^2(\Omega) \cap H^1_0(\Omega)$ \cite{Tho}.

Through this paper, we denote $\alpha_0:= \alpha(0)$ and assume that $\alpha(t)$ has bounded second-order derivative on $[0, T]$.  In addition, we use $Q$ to denote a generic positive constant that may assume different values at different occurrences. For simplicity, we may drop the domain $\om$ and the time interval $(0,T)$ in the Sobolev spaces and norms, e.g. we write $\| \cdot\|_{L^2(L^2)}$ instead of $\| \cdot\|_{L^2(0,T;L^2(\Omega))}$, when no confusion occurs.

\subsection{Spectral decomposition}
For $t \in [0, T]$, we expand $u$ and $f$ with respect to  $\{\phi_i\}_{i=1}^{\infty}$ as follows \cite{SakYam}
$$u=\sum_{i=1}^\infty u_i(t)\phi_i(\bm x),~u_i(t) := \big (u(\cdot,t),\phi_i \big),~~f=\sum_{i=1}^\infty f_i(t)\phi_i(\bm x),~f_i(t) := \big (f(\cdot,t),\phi_i \big).$$
We invoke these expressions into (\ref{VtFDEs}) to find that $\{u_i\}_{i=1}^{\infty}$ satisfy the following integro-differential equations
\begin{align}\label{ode1}
u_i'+\lambda_i k*u_i=f_i,\text{ for }t\in (0,T];~~u_i(0)=u_{0,i}:=(u_0,\phi_i), \quad i=1,2, \cdots
\end{align}

 For $\mu>0$, we first define
$\beta_\mu (t):=\frac{t^{\mu-1}}{\Gamma(\mu)}$.  We follow the idea of the perturbation method proposed in \cite{Zhe} to split the kernel $k(t)$ as
\begin{align}
k(t)&=\frac{t^{-\alpha(t)}}{\Gamma(1-\alpha(t))}=\frac{t^{-\alpha_0}}{\Gamma(1-\alpha_0)}+\int_0^t\p_z\frac{t^{-\alpha(z)}}{\Gamma(1-\alpha(z))}dz=:\beta_{1-\alpha_0}+\tilde g(t).\label{hh3}
\end{align}
It is shown in \cite{Zhe} that
\begin{align}\label{dbndg}
|\tilde g|\leq Q\int_0^t t^{-\alpha_0}(1+|\ln t|)dz\leq Qt^{1-\alpha_0}(1+|\ln t|),~~|\tilde g'|\leq Qt^{-\alpha_0}(1+|\ln t|).
\end{align}
We thus employ (\ref{hh3}) to rewrite (\ref{ode1}) as
\begin{align}\label{ode2}
u_i'+\lambda_i \beta_{1-\alpha_0}*u_i=f_i-\lambda_i\tilde g*u_i,\text{ for }t\in (0,T];~~u_i(0)=u_{0,i}.
\end{align}

\section{Mathematical analysis}\label{sec3}

We prove the well-posedness and the regularity results  of (\ref{VtFDEs}).

\begin{theorem}\label{thm1:pde}
Suppose  $f \in H^1(L^2)$ and $u_{0} \in \check H^2$, the problem (\ref{VtFDEs})  has a unique solution $u \in H^1(L^2)$ and
\begin{equation}\label{pdestab1}\begin{array}{l}
\ds \|u\|_{H^1(L^2)}\leq Q\big(\|u_{0}\|_{\check H^2}+ \|f\|_{H^1(L^2)}\big).
\end{array}
\end{equation}
\end{theorem}
\textbf{Proof}
The proof of the theorem  will be carried out in two steps.

\textit{Step A: Well-posedness of problem (\ref{ode2}).}

We first consider the case $u_{0,i}=0$ in (\ref{ode2}) and let $\mathcal X := \{ q \in H^1(0,T): q(0)=0 \}$ equipped with the equivalent norm $\|q\|_{\mathcal X,\sigma} := \|e^{-\sigma t} q'\|_{L^2(0,T)}$ for some $\sigma\geq 0$. For each $v\in \mathcal X$, let $w:= \mathcal M v$ be the solution of
\begin{align}\label{ode4}
w'+\lambda_i \beta_{1-\alpha_0}*w=f_i-\lambda_i\tilde g*v,\text{ for }t\in (0,T];~~w(0)=0.
\end{align}
Then  the solution $w$ to (\ref{ode4}) could be represented as \cite{Mcl}
\begin{align}\label{OdeSol}
w=E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0})*(f_i-\lambda_i\tilde g*v).
\end{align}
To bound $\|w\|_{\mathcal X,\sigma}$, we directly differentiate the above equation to get
\begin{align}\label{w'}
w'=E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0})f_i(0)+E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0})*(f_i'-\lambda_i\tilde g'*v).
\end{align}
By using the Mittag-Leffler function \cite{DieFor,Gor,Pod} and the fact that $$[tE_{2-\alpha_0,2}(-\lambda_it^{2-\alpha_0})\big]_t = E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0}),$$ we further apply integration by parts to reformulate the last term on the right-hand side of (\ref{w'}) to get
\begin{equation}\label{ode5}\begin{array}{rl}
\hspace{-0.1in} \ds E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0})*(\lambda_i\tilde g'*v)&\ds =tE_{2-\alpha_0,2}(-\lambda_it^{2-\alpha_0})*(\lambda_i\tilde g'*v')\\[0.1in]
\hspace{-0.1in} \ds &\ds =[t^{\alpha_0-1}(\lambda_i t^{2-\alpha_0})E_{2-\alpha_0,2}(-\lambda_it^{2-\alpha_0})]*(\tilde g'*v').
\end{array}
\end{equation}
By the asymptotic property of the Mittag-Leffler function \cite{Gor,Jinbook},  we have
\begin{align}\label{MLbound}
\big|(\lambda_i t^{2-\alpha_0})E_{2-\alpha_0,2}(-\lambda_it^{2-\alpha_0})\big|\leq Q,
\end{align}
where $Q$ is independent from $t$ and $\lambda_i$. Thus we apply Young's convolution inequality and the estimate
\begin{equation}\label{AdjEq:e4}
\int_0^Te^{-\sigma t }t^{\alpha_0-1}dt\leq \sigma^{-\alpha_0}\int_0^{\infty}e^{-z}z^{\alpha_0-1}dz\leq Q\sigma^{-\alpha_0},
\end{equation}
and (\ref{MLbound}) to bound (\ref{ode5}) as
\begin{equation*}\begin{array}{rl}
&\big\|e^{-\sigma t}[E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0})*(\lambda_i\tilde g'*v)]\big\|_{L^2(0,T)}\\[0.1in]
& \qquad\ds\leq Q\big\|e^{-\sigma t}(t^{\alpha_0-1}*|\tilde g'*v'|)\big\|_{L^2(0,T)}\\[0.1in]
&\qquad\ds\leq Q\big\|(e^{-\sigma t}t^{\alpha_0-1})*|e^{-\sigma t}\tilde g'|*|e^{-\sigma t}v'|\big\|_{L^2(0,T)}\\[0.1in]
&\qquad \ds\leq Q\|e^{-\sigma t}t^{\alpha_0-1}\|_{L^1(0,T)}\|e^{-\sigma t}\tilde g'\|_{L^1(0,T)}\|v\|_{\mathcal X,\sigma}\leq Q\sigma^{-\alpha_0}\|v\|_{\mathcal X,\sigma}.
\end{array}
\end{equation*}
We invoke this estimate, the boundedness of $E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0})$ \cite{Gor,Jinbook} and $|f_i(0)|\leq Q\|f_i\|_{H^1(0,T)}$ in (\ref{w'}) to get
\begin{align}\label{ode7}
\|w\|_{\mathcal X,\sigma}\leq Q\|f_i\|_{H^1(0,T)}+Q\sigma^{-\alpha_0}\|v\|_{\mathcal X,\sigma}.
\end{align}
Thus the mapping $\mathcal M:\mathcal X\rightarrow \mathcal X$ is well defined. To show its contractivity, let $w_i=\mathcal M v_i$ for $i =1$, $2$,  $e_w : = w_1- w_2$ and $e_v : = v_1- v_2$ satisfy
\begin{align}\label{ode6}
e_w'+\lambda_i \beta_{1-\alpha_0}*e_w=-\lambda_i\tilde g*e_v.
\end{align}
 Then an application of the estimate (\ref{ode7}) to (\ref{ode6})  gives
 \begin{align}\label{ode8}
\|e_w\|_{\mathcal X,\sigma}\leq Q\sigma^{-\alpha_0}\|e_v\|_{\mathcal X,\sigma}.
\end{align}
 We select a sufficiently large $\sigma$ such that the mapping $\mathcal M$ is a contraction. By the Banach fixed point theorem, $\mathcal M$ has a unique fixed point $v=\mathcal Mv$ with the estimate $\|v\|_{\mathcal X,\sigma}\leq Q\|f_i\|_{H^1(0,T)}$.

For the case that $u_{0,i}\neq 0$, we could directly verify that
$$u_i=E_{2-\alpha_0,1}(-\lambda_it^{2-\alpha_0})u_{0,i}+v,$$
where $v$ is the fixed point of $\mathcal M$, solves
(\ref{ode2}) and thus (\ref{ode1}). Differentiate this equation to get
$$u_i'=-\lambda_i t^{1-\alpha_0}E_{2-\alpha_0,2-\alpha_0}(-\lambda_it^{2-\alpha_0})u_{0,i}+v',$$
which, together with $\|v\|_{\mathcal X,\sigma}\leq Q\|f_i\|_{H^1(0,T)}$, leads to
\begin{equation}\label{odestab}\begin{array}{l}
\ds \|u_i\|_{H^1(0,T)}\leq Q\big(\lambda_i|u_{0,i}|+\|f_i\|_{H^1(0,T)}\big).
\end{array}
\end{equation}
The uniqueness of (\ref{ode1}) follows from that of (\ref{ode2}) with $u_{0,i}=0$. Consequently, we conclude that  (\ref{ode2}) and thus (\ref{ode1}) admit  a unique solution in $H^1(0,T)$ with the estimate (\ref{odestab}).

\textit{Step B: Well-posedness of the problem  (\ref{VtFDEs}).}

We invoke the  estimate   (\ref{odestab}) to bound
 $   \bar u:=\int_0^t\big[\sum_{i=1}^\infty u_i'(s)\phi_i(\bm x)\big]ds + u_0
$
by using
\begin{equation}\begin{array}{rl}\label{AdjEq:e8}
\ds \|\bar u\|_{H^1(L^2)}^2 \leq Q \|\p_t \bar u\|^2_{L^2(L^2)} = \sum_{i=1}^\infty \|u_i'\|_{L^2(0,T)}^2
&\hspace{-0.1in}\ds \leq  Q\sum_{i=1}^\infty \big(\lambda_i^2|u_{0,i}|^2 + \|f_i\|_{H^1(0,T)}^2\big)\\[0.15in]
&\hspace{-0.1in}\ds  =Q\big(\|u_0\|_{\check H^2}^2 + \|f\|^2_{H^1(L^2)}\big).
\end{array}
\end{equation}
As$u_i(t) = \int_0^t u_i'(s)ds + u_{0,i}$ satisfies the differential equation in (\ref{ode1}) for $i \ge 1$, $\bar u \in H^1(L^2)$ is a solution to problem (\ref{VtFDEs}).
The uniqueness of the solution to the problem (\ref{VtFDEs})   follows from that of the differential equation in (\ref{ode1}). We thus complete the proof of the theorem.

\begin{theorem}\label{thm:pde}
Suppose  $f \in H^1(\check H^2)$ and $u_{0} \in \check H^4$, the following regularity estimate holds for the  problem \eqref{VtFDEs}
\begin{equation}\label{pde:reg}\begin{array}{l}
\ds  \big\|t^{\alpha_0/2} \p_t^2 u \big\|_{L^2(L^2)}  +  \| u  \|_{H^1(\check H^2)} \le Q\big(\|u_{0}\|_{\check H^4}+ \|f\|_{H^1(\check H^2)}\big).
\end{array}
\end{equation}
\end{theorem}
\textbf{Proof}
We  differentiate the equation in (\ref{ode2}) to obtain
 \begin{align}\label{odeutt}
u_i^{\prime \prime}+\lambda_i \beta_{1-\alpha_0}*u_i^{\prime}=f_i^{\prime}-\lambda_i\tilde g^{\prime}*u_i-\lambda_i \beta_{1-\alpha_0}u_{0,i},
\end{align}
which, combined with the first estimate in (\ref{odestab}), (\ref{dbndg}) and Young’s convolution
inequality, gives
\begin{equation}\label{utt}\begin{array}{rl}
\ds \big\|t^{\alpha_0/2} u_i^{\prime \prime}\big\|_{L^2(0, T)} &\hspace{-0.1in}\ds  \le Q \big(\lambda_i \|\beta_{1-\alpha_0}\|_{L^1(0, T)}\|u_i^{\prime}\|_{L^2(0, T)} + \|f_i^{\prime}\|_{L^2(0, T)} \\[0.1in]
&\hspace{-0.1in}\ds  \qquad + \lambda_i \|\tilde g^{\prime}\|_{L^1(0, T)}\|u_i\|_{L^2(0, T)} + \lambda_i|u_{0,i}| \big)\\[0.1in]
&\hspace{-0.1in}\ds \le Q\big(\lambda_i\|u_i\|_{H^1(0, T)} + \|f_i^{\prime}\|_{L^2(0, T)} + \lambda_i|u_{0,i}|  \big)\\[0.1in]
&\hspace{-0.1in}\ds   \le Q\big(\lambda_i^2|u_{0,i}|+ \lambda_i \|f_i\|_{H^1(0,T)}\big).
\end{array}
\end{equation}
We further combine   (\ref{odestab}) and (\ref{utt}) and follow the procedures in (\ref{AdjEq:e8})  to obtain
\begin{equation*}\begin{array}{l}
\ds \big \|t^{{\alpha_0}/{2}}\p_t^2 u \big \|^2_{L^2(L^2)} +  \| u  \|^2_{H^1(\check H^2)}  = \sum_{i=1}^\infty \big( \|t^{{\alpha_0}/{2}} u_i^{\prime \prime}\|_{L^2(0,T)}^2 + \lambda_i^2\| u_i\|_{H^1(0,T)}^2 \big)\\[0.1in]
\qquad \qquad \ds \leq Q\sum_{i=1}^\infty \big( \lambda_i^2\|f_i\|^2_{H^1(0,T)} + \lambda_i^4 |u_{0,i}|^2 \big)\le Q \big ( \|f\|^2_{H^1(\check H^2)} + \|u_0\|^2_{\check H^4} \big),
\end{array}\end{equation*}
which gives (\ref{pde:reg}) and we thus complete the proof.

\section{Concluding remarks}

This work establishes the well-posedness and regularity results of an evolution equation with non-positive type variable-exponent memory.
A perturbation method is employed to reformulate the model, based on which the analysis is performed. Based on these theoretical results, we may consider error estimates of the numerical approximation to model \eqref{VtFDEs}. However, the coupling of the variable-exponent kernel and the Laplacian may cause difficulties in  stability and error estimates that have not been encountered before. We will investigate this interesting topic in the near future.

\section*{Acknowledgments}
This work was partially supported by the National Key R\&D Program of China (No. 2023YFA1008903), the National Natural Science Foundation of China (No. 12301555), the Taishan Scholars Program of Shandong Province (No. tsqn202 306083),  the China Postdoctoral Science Foundation (No. 2024M762459) and the Natural Science Foundation of Hubei Province (No. JCZRQN202500278).

\end{document}